# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v4

@@ II-C Vision Transformer Token Compression
III Image-centric Token Compression
Multimodal long context token compression methods generally fall into four categories based on their underlying mechanisms: transformation-based approaches directly transform the cross-modality information to compress tokens by altering their scale or representation; similarity-based techniques reduce tokens by leveraging the inherent resemblances between them; attention-based strategies exploit the sparsity of attention within the multimodal data to guide compression; and query-based methods selectively refine multimodal information, guided by prompts, to distill the most relevant tokens. Each of these methods has its own set of advantages and disadvantages, which are summarized in Table I. Representative image-centric token compression methods are further compared in Table II.
III-A Transformation-based Image-centric Compression
@@
III-A1 Pixel Unshuffle
Pixel unshuffle is the inverse operation of pixel shuffle. It transforms a feature map from a high spatial resolution with a small number of channels into a lower-resolution feature map with a larger number of channels. This effectively reduces the number of tokens. The transformation can be mathematically expressed as:
@@ III-A1 Pixel Unshuffle
where $H$, $W$ denote the height and width of the token grid, $D$ is the hidden dimension of each token, and $r$ is the downsampling ratio.
Recent works like InternVL series [30, 158, 159, 160], Qwen2 series [32, 33], and NVLM [31] utilize pixel unshuffle to reduce the number of tokens generated by the vision tower by a factor of one-quarter. Subsequently, an MLP is employed to align the visual dimension with the text dimension, addressing the mismatch in the hidden dimension.
III-A2 Spatial Pooling / Interpolation
@@
LLaVA-OneVision [2] employs bilinear interpolation for 2D downsampling of aligned tokens, while LLaVA-Video [34] uses average pooling for downsampling. M3 [35] utilizes a simple pooling operation to learn an inherently multi-granular representation during training. This allows the model to achieve comparable performance with fewer tokens during inference, effectively addressing efficiency concerns. DeCo [36] argues that the Q-former [1, 81] is an inefficient visual compressor and similarly achieves token compression through a straightforward average pooling approach, leading to improved convergence efficiency and performance.
III-A3 Spatial Convolution
@@
Honeybee [37] proposes the C-Abstractor, which uses convolution to extract and compress token information while preserving locality. MobileVLM [38], on the other hand, employs an LDP module that utilizes depth-wise convolution to reduce the number of tokens by $75\%$.
III-A4 Comparative Analysis of Transformation Methods
@@
III-B Similarity-based Image-centric Compression
@@
Early works in this area include ToMe [60], an acceleration method for ViTs. ToMe introduces a token merge module between the attention and MLP blocks, calculating token similarity and merging similar tokens via bipartite soft matching. This process creates a new set of tokens, $\mathcal{T}$, by replacing the most similar tokens with their merged representations.
@@
In the context of MLLMs, FOLDER [61] employs a similar approach, inserting a token merge module within the last attention block of the vision encoder. This reduces the number of tokens that were subsequently passed to the LLM decoder. DivPrune [28] reframes the token compression problem as a Max-Min diversity problem [161], aiming to select a subset of tokens with maximal internal differences. AuroraCap [45] adopts a strategy consistent with ToMe, performing token merging within each attention and MLP block of the vision tower. This progressively reduces the number of tokens throughout the ViT model. While the aforementioned methods primarily leverage similarity-based clustering of tokens within the ViT, TopV [49] extends this principle to compress tokens within the LLM layers. TopV comprehensively considers both the similarity and distance functions between features to guide the token compression process, operating directly within the multimodal representation space of the LLM.
III-C Attention-based Image-centric Compression
Attention-based token compression methods leverage the inherent sparsity of visual feature attention to guide token pruning. Tokens with low attention scores can often be considered removable without significantly impacting the original computation. In vision language models, both the vision encoder and the LLM decoder incorporate transformers. Consequently, attention-based compression strategies can be broadly categorized into those applied within the encoder and those within the decoder.
TABLE II: Comparison of Training-Free Token Compression Methods for image LLMs in Image Understanding Tasks
@@ III-C Attention-based Image-centric Compression
BLIP-2 [81] 32 224 65.0 41.0 19.6 61.0 42.5 85.3 1293.8 46.4 38.1 22.4
IDEFICS-9B [162] 64 224 50.9 38.4 35.5 25.9 48.2
MobileVLM-3B [39] 144 336 59.0 61.0 47.5 84.9 1288.9 59.6
mPLUG-Owl2 [82] 1024 448 79.4 56.1 54.5 68.7 54.3 1450.2 64.5 57.8 36.2
Video-LLaVA [6] 256 224 74.7 60.3 48.1 66.4 51.8 84.4 60.9 73.1 32.0
Qwen-VL [32] 256 448 78.8 59.3 35.2 67.1 63.8 38.2 56.3
LLaVA-v1.5 [1] 576 336 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.6 63.4 30.5
@@
ToMe [60] 192 336 68.0 54.3 52.1 1563.0 60.5
FastV [26] 192 336 67.1 52.7 52.5 64.8 1612.0 61.2 57.1 27.7
SparseVLM [51] 192 336 75.6 57.6 56.1 83.6 1721.0 62.5 55.8 31.5
MustDrop [52] 192 336 76.0 58.2 51.4 56.5 1787.0 62.3
PruMerge+ [29] 144 336 76.8 68.3 57.1 84.0 1462.4 64.9
ATP-LLaVA [76] 144 336 76.4 59.5 84.2 1473.9 66.0 57.3 31.5
VisionZip++ [18] 128 336 76.6 58.9 57.0 83.7 1823.0 55.8 32.9
VisPruner [65] 128 336 75.8 58.2 52.7 57.0 84.6 1461.4 62.7 33.7
VisionZip++ [18] 64 336 74.2 57.0 56.0 80.9 1756.0 53.4 30.2
TokenCarve [78] 64 336 74.8 57.0 79.9 1754.0 62.0 29.3
VisPruner [65] 32 336 67.7 52.2 53.0 53.9 72.7 1271.0 58.4 28.8
@@
LLaMA-VID [40] 2 336 55.5 68.8 49.0 83.1
LLaVA-Mini [85] 1 336 77.6 60.9 56.2 70.4 57.0 84.7 1466.0 65.6 58.5 68.9 36.6
VoCo-LLAMA [86] 1 336 72.3 57.0 65.4 81.4 1323.3 58.8 53.7
III-C1 Attention in Encoder
@@
Prumerge [29] selects cluster centers for visual tokens based on [CLS] attention in the encoder. It then merges the remaining less attentive tokens using K-nearest neighbors (KNN) clustering and a weighted cluster center update mechanism. VisionZip [18] retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. VisPruner [65] similarly preserves a subset of high-attention visual tokens. Then it progressively removes duplicates based on similarity in multiple rounds, ultimately retaining an additional set of diverse tokens. GlobalCom2 [69] employs a hierarchical strategy. It coordinates the attention scores of thumbnail tokens to guide the pruning of high-resolution crops, thereby achieving effective global context reduction.
III-C2 Attention in Decoder
@@
FastV [26] is among the first to identify a significant inefficiency in large vision language models (LVLMs), namely the extremely low attention efficiency of visual tokens. For instance, in LLaVA-v1.5, visual tokens received only $0.21\%$ of the attention obtained by system prompts after the second layer. FastV posits that this is due to an overabundance of visual signals, leading to specific features aggregating onto “anchor” tokens via shallow self-attention mechanisms. Consequently, pruning $50\%$ of visual tokens based on attention scores after the second layer maintains maximal performance. PyramidDrop [72] structures the token compression process within the LLM into multiple stages. It employs progressive token compression to avoid excessive loss of visual information in shallower layers. VTW [73] takes a more aggressive pruning approach, arguing that visual tokens can be entirely removed after a certain layer within the LLM. The specific layer for visual token removal is determined using a calibration dataset. FitPrune [74] focuses on reducing the length of visual tokens per layer. It considers both the self-attention of visual tokens and their cross-attention with text tokens to guide compression. The goal is to find an optimal pruning “recipe” that minimizes the distributional gap before and after pruning. ST3 [75] dynamically reduces tokens during the generation process. It also progressively prunes inattentive visual tokens as the layer goes deeper. ATP-LLaVA [76] introduces an adaptive token pruning (ATP) module within the decoder layers. This module trains threshold heads to adaptively predict pruning thresholds for the current layer and instance, thereby removing redundant or text-irrelevant visual tokens. ZipVL [46] achieves progressive compression by determining the compression ratio for each layer based on its attention score distribution. This allows for a granular and adaptive reduction of visual tokens throughout the model.
III-C3 Critical Challenge for Pruning in Decoder
While these methods leverage attention scores within the LLM decoder to offer sophisticated ways to compress visual tokens, they face a significant practical challenge: the explicit need to access attention scores. This direct access is often incompatible with highly optimized acceleration libraries like FlashAttention [163, 164], which compute attention implicitly or in a fused manner for speed. This incompatibility can be mitigated by performing an additional, separate attention calculation solely for pruning purposes. However, for progressive pruning strategies such as FitPrune, ST3, and ZipVL, this additional computational overhead becomes significantly more pronounced, potentially negating the efficiency gains.
III-D Query-based Image-centric Compression
@@
III-D1 Token Distillation
Token distillation originates from the early projector designs of MLLM. The goal is to distill visual tokens to learn the most text-relevant visual representations, reduce visual tokens while simultaneously align modalities.
The Q-Former series [1, 81], a pioneering approach, uses learnable queries and cross-attention to extract pertinent visual cues from visual features. Similarly, mPLUG-Owl [82], MiniGPT-4 [83], Flamingo [84], and Qwen-VL [4] all employ variations of learnable query-based architectures to condense visual information into a smaller fixed set of tokens that are then aligned with the language model. LLaMA-VID [40] employs a highly aggressive approach to visual token compression. For a single image or video frame, it utilizes context attention where the text query aggregates text-related visual cues from the visual embedding. Ultimately, it represents an entire image’s information using only two tokens. LLaVA-Mini [85] achieves comparable performance by pre-fusing visual information directly into text tokens, requiring just one visual token. While previous methods relied on external modules for visual token compression, VoCo-LLaMA [86] is notable as the first approach to use LLMs themselves for visual token compression. It distills the LLM’s understanding of visual tokens into the processing of VoCo tokens via attention distillation. Victor [87] introduces a small number of learnable “register tokens” after the visual tokens. It then uses the shallow layers of a large model to distill visual information into these registers, discarding all original visual tokens to significantly improve inference and training efficiency.
III-D2 Cross-Modal Selection
@@
SparseVLM [51] employs visual tokens to pre-select relevant text tokens. By leveraging the visual modality as an initial filter, SparseVLM efficiently narrows down the textual search space, focusing on information pertinent to the visual content. AdaFV [88] employs a dual-metric approach for selecting the most informative visual tokens. It calculates both text-to-image similarity and visual saliency extracted from the vision encoder. By combining these two indicators, AdaFV identifies visual tokens that are not only semantically aligned with the text but also visually prominent or significant. TRIM [89] introduces a unique method that begins by identifying outlier tokens based on the similarity between text and visual tokens; these outliers are deemed important. Subsequently, a clustering algorithm is utilized to merge the remaining, less critical tokens. This approach prioritizes distinct, highly relevant tokens before consolidating the rest.
TABLE III: Comparison of Training-Free Token Compression Methods for video LLMs in Video Understanding Tasks

### v5

@@ 2.4 Problem Definition and Taxonomy Scope
3 Image-centric Token Compression
Multimodal long context token compression methods generally fall into four categories based on their underlying mechanisms: transformation-based approaches directly transform the cross-modality information to compress tokens by altering their scale or representation; similarity-based techniques reduce tokens by leveraging the inherent resemblances between them; attention-based strategies exploit the sparsity of attention within the multimodal data to guide compression; and query-based methods selectively refine multimodal information, guided by prompts, to distill the most relevant tokens. Each of these methods has its own set of advantages and disadvantages, which are summarized in Table 1. Representative image-centric token compression methods are further compared in Table 2.
3.1 Transformation-based Image-centric Compression
@@
3.1.1 Pixel Unshuffle
Pixel unshuffle is the inverse operation of pixel shuffle. It transforms a feature map from a high spatial resolution with a small number of channels into a lower-resolution feature map with a larger number of channels. This reduces the number of tokens. The transformation can be mathematically expressed as:
@@ 3.1.1 Pixel Unshuffle
where $H$, $W$ denote the height and width of the token grid, $D$ is the hidden dimension of each token, and $r$ is the downsampling ratio. Here $r$ is a positive integer, typically 2. Therefore, as summarized in Table 1, the token compression ratio for transformation-based methods is usually limited to a few specific values, generally compressing the number of tokens to 25%.
Recent works like InternVL series (Chen et al., 2024c; b; Gao et al., 2024b; Zhu et al., 2025a), Qwen2 series (Wang et al., 2024b; Bai et al., 2025), and NVLM (Dai et al., 2024) utilize pixel unshuffle to reduce the tokens generated by the vision tower by a factor of one-quarter. Subsequently, an MLP is employed to align the visual dimension with the text dimension, addressing the mismatch in the hidden dimension.
3.1.2 Spatial Pooling / Interpolation
@@
LLaVA-OneVision (Li et al., 2025a) employs bilinear interpolation for 2D downsampling of aligned tokens, while LLaVA-Video (Zhang et al., 2024d) uses average pooling for downsampling. M3 (Cai et al., 2024a) utilizes a simple pooling operation to learn an inherently multi-granular representation during training. This allows the model to achieve comparable performance with fewer tokens during inference, effectively addressing efficiency concerns. DeCo (Yao et al., 2024) argues that the Q-former (Liu et al., 2023; Li et al., 2023c) is an inefficient visual compressor and similarly achieves token compression through a straightforward average pooling approach, leading to improved convergence efficiency and performance.
3.1.3 Spatial Convolution
@@
Honeybee (Cha et al., 2024) proposes the C-Abstractor, which uses convolution to extract and compress token information while preserving locality. MobileVLM (Chu et al., 2023), on the other hand, employs an LDP module that utilizes depth-wise convolution to reduce the number of tokens by $75\%$.
3.1.4 Comparative Analysis of Transformation Methods
@@
3.2 Similarity-based Image-centric Compression
@@
Early works in this area include ToMe (Bolya et al., 2023), an acceleration method for ViTs. ToMe introduces a token merge module between the attention and MLP blocks, calculating token similarity and merging similar tokens via bipartite soft matching. This process creates a new set of tokens, $\mathcal{T}$, by replacing the most similar tokens with their merged representations.
@@
In the context of MLLMs, FOLDER (Wang et al., 2025a) employs a similar approach, inserting a token merge module within the last attention block of the vision encoder. This reduces the number of tokens that were subsequently passed to the LLM decoder. DivPrune (Alvar et al., 2025) reframes the token compression problem as a Max-Min diversity problem (Porumbel et al., 2011), aiming to select a subset of tokens with maximal internal differences. AuroraCap (Chai et al., 2025) adopts a strategy consistent with ToMe, performing token merging within each attention and MLP block of the vision tower. This progressively reduces the number of tokens throughout the ViT model. While the aforementioned methods primarily leverage similarity-based clustering of tokens within the ViT, TopV (Yang et al., 2025a) extends this principle to compress tokens within the LLM layers. TopV comprehensively considers both the similarity and distance functions between features to guide the token compression process, operating directly within the multimodal representation space of the LLM.
3.2.1 Analysis of Similarity Methods
While similarity-based methods effectively reduce tokens, this merging often overlooks the original spatial information of the tokens, leading to spatial misunderstanding (in Tab. 1). Subsequent work frequently employs methods like DPC-KNN (Du et al., 2016; Rodriguez and Laio, 2014) or techniques focused on local spatial similarity merging to prevent excessive spatial information degradation. Furthermore, when tokens are over-generalized, similarity-based methods struggle to distinguish between them, easily leading to misjudgment.
3.3 Attention-based Image-centric Compression
Attention-based token compression methods leverage the inherent sparsity of visual feature attention to guide token pruning. Tokens with low attention scores can often be considered removable without significantly impacting the original computation. Specifically, these methods utilize the attention mechanism to identify and preserve pivotal tokens. It is worth noting that this shares the same underlying philosophy as sparse attention methods (Yuan et al., 2025; Zhang et al., 2025c; Lu et al., 2025; Yin et al., 2025), which focus on executing the critical attention computations, yet manifest at a different scale: the former operates on token quantity while the latter operates on computational pathways. In vision language models, both the vision encoder and the LLM decoder incorporate transformers. Consequently, attention-based compression strategies can be broadly categorized into those applied within the encoder and those within the decoder.
Table 2: Comparison of Training-Free Token Compression Methods for Image LLMs in Understanding Tasks
@@ 3.3 Attention-based Image-centric Compression
BLIP-2 (Li et al., 2023c) 32 224 65.0 41.0 19.6 61.0 42.5 85.3 1293.8 46.4 38.1 22.4
IDEFICS-9B (Laurençon et al., 2023) 64 224 50.9 38.4 35.5 25.9 48.2
MobileVLM-3B (Chu et al., 2024a) 144 336 59.0 61.0 47.5 84.9 1288.9 59.6
mPLUG-Owl2 (Ye et al., 2023) 1024 448 79.4 56.1 54.5 68.7 54.3 1450.2 64.5 57.8 36.2
Video-LLaVA (Lin et al., 2024a) 256 224 74.7 60.3 48.1 66.4 51.8 84.4 60.9 73.1 32.0
Qwen-VL (Wang et al., 2024b) 256 448 78.8 59.3 35.2 67.1 63.8 38.2 56.3
LLaVA-v1.5 (Liu et al., 2023) 576 336 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.6 63.4 30.5
@@
ToMe (Bolya et al., 2023) 192 336 68.0 54.3 52.1 1563.0 60.5
FastV (Chen et al., 2024a) 192 336 67.1 52.7 52.5 64.8 1612.0 61.2 57.1 27.7
SparseVLM (Zhang et al., 2024c) 192 336 75.6 57.6 56.1 83.6 1721.0 62.5 55.8 31.5
MustDrop (Liu et al., 2024c) 192 336 76.0 58.2 51.4 56.5 1787.0 62.3
PruMerge+ (Shang et al., 2025) 144 336 76.8 68.3 57.1 84.0 1462.4 64.9
ATP-LLaVA (Ye et al., 2025b) 144 336 76.4 59.5 84.2 1473.9 66.0 57.3 31.5
VisionZip++ (Yang et al., 2025c) 128 336 76.6 58.9 57.0 83.7 1823.0 55.8 32.9
VisPruner (Zhang et al., 2025e) 128 336 75.8 58.2 52.7 57.0 84.6 1461.4 62.7 33.7
VisionZip++ (Yang et al., 2025c) 64 336 74.2 57.0 56.0 80.9 1756.0 53.4 30.2
TokenCarve (Tan et al., 2025) 64 336 74.8 57.0 79.9 1754.0 62.0 29.3
VisPruner (Zhang et al., 2025e) 32 336 67.7 52.2 53.0 53.9 72.7 1271.0 58.4 28.8
@@
LLaMA-VID (Li et al., 2024c) 2 336 55.5 68.8 49.0 83.1
LLaVA-Mini (Zhang et al., 2025f) 1 336 77.6 60.9 56.2 70.4 57.0 84.7 1466.0 65.6 58.5 68.9 36.6
VoCo-LLAMA (Ye et al., 2025c) 1 336 72.3 57.0 65.4 81.4 1323.3 58.8 53.7
3.3.1 Attention in Encoder
@@
Prumerge (Shang et al., 2025) selects cluster centers for visual tokens based on [CLS] attention in the encoder. It then merges the remaining less attentive tokens using K-nearest neighbors (KNN) clustering and a weighted cluster center update mechanism. VisionZip (Yang et al., 2025c) retains visual tokens with high attention scores and subsequently merges the remaining tokens through clustering. VisPruner (Zhang et al., 2025e) similarly preserves a subset of high-attention visual tokens. Then it progressively removes duplicates based on similarity in multiple rounds, ultimately retaining an additional set of diverse tokens. GlobalCom2 (Liu et al., 2025a) employs a hierarchical strategy. It coordinates the attention scores of thumbnail tokens to guide the pruning of high-resolution crops, thereby achieving effective global context reduction.
3.3.2 Attention in Decoder
@@
FastV (Chen et al., 2024a) is among the first to identify a significant inefficiency in large vision language models (LVLMs), namely the extremely low attention efficiency of visual tokens. For instance, in LLaVA-v1.5, visual tokens received only $0.21\%$ of the attention obtained by system prompts after the second layer. FastV posits that this is due to an overabundance of visual signals, leading to specific features aggregating onto “anchor” tokens via shallow self-attention mechanisms. Consequently, pruning $50\%$ of visual tokens based on attention scores after the second layer maintains maximal performance. PyramidDrop (Xing et al., 2025) structures the token compression process within the LLM into multiple stages. It employs progressive token compression to avoid excessive loss of visual information in shallower layers. VTW (Lin et al., 2025c) takes a more aggressive pruning approach, arguing that visual tokens can be entirely removed after a certain layer within the LLM. The specific layer for visual token removal is determined using a calibration dataset. FitPrune (Ye et al., 2025a) focuses on reducing the length of visual tokens per layer. It considers both the self-attention of visual tokens and their cross-attention with text tokens to guide compression. The goal is to find an optimal pruning “recipe” that minimizes the distributional gap before and after pruning. ST3 (Zhuang et al., 2025) dynamically reduces tokens during the generation process. It also progressively prunes inattentive visual tokens as the layer goes deeper. ATP-LLaVA (Ye et al., 2025b) introduces an adaptive token pruning (ATP) module within the decoder layers. This module trains threshold heads to adaptively predict pruning thresholds for the current layer and instance, thereby removing redundant or text-irrelevant visual tokens. ZipVL (He et al., 2025) achieves progressive compression by determining the compression ratio for each layer based on its attention score distribution. This allows for a granular and adaptive reduction of visual tokens throughout the model.
3.3.3 Critical Challenge for Pruning in Decoder
While these methods leverage attention scores within the LLM decoder to offer sophisticated ways to compress visual tokens, they face a significant practical challenge: the explicit need to access attention scores. This direct access is often incompatible with highly optimized acceleration libraries like FlashAttention (Dao et al., 2022; Dao, 2024), which compute attention implicitly or in a fused manner for speed. This incompatibility can be mitigated by performing an additional, separate attention calculation solely for pruning purposes. However, for progressive pruning strategies such as FitPrune, ST3, and ZipVL, this additional computational overhead becomes significantly more pronounced, potentially negating the efficiency gains.
3.4 Query-based Image-centric Compression
@@
3.4.1 Token Distillation
Token distillation originates from the early projector designs of MLLM. The goal is to distill visual tokens to learn the most text-relevant visual representations, reduce visual tokens while also aligning modalities.
The Q-Former series (Liu et al., 2023; Li et al., 2023c), a pioneering approach, uses learnable queries and cross-attention to extract pertinent visual cues from visual features. Similarly, mPLUG-Owl (Ye et al., 2023), MiniGPT-4 (Zhu et al., 2024), Flamingo (Alayrac et al., 2022), and Qwen-VL (Bai et al., 2023) all employ variations of learnable query-based architectures to condense visual information into a smaller fixed set of tokens that are then aligned with the language model. LLaMA-VID (Li et al., 2024c) employs a highly aggressive approach to visual token compression. For a single image or video frame, it utilizes context attention where the text query aggregates text-related visual cues from the visual embedding. Ultimately, it represents an entire image’s information using only two tokens. LLaVA-Mini (Zhang et al., 2025f) achieves comparable performance by pre-fusing visual information directly into text tokens, requiring just one visual token. While previous methods relied on external modules for visual token compression, VoCo-LLaMA (Ye et al., 2025c) is notable as the first approach to use LLMs themselves for visual token compression. It distills the LLM’s understanding of visual tokens into the processing of VoCo tokens via attention distillation. Victor (Wen et al., 2024) introduces a small number of learnable “register tokens” after the visual tokens. It then uses the shallow layers of a large model to distill visual information into these registers, discarding all original visual tokens to significantly improve inference and training efficiency.
3.4.2 Cross-Modal Selection
@@
SparseVLM (Zhang et al., 2024c) employs visual tokens to pre-select relevant text tokens. By leveraging the visual modality as an initial filter, SparseVLM efficiently narrows down the textual search space, focusing on information pertinent to the visual content. AdaFV (Han et al., 2025) employs a dual-metric approach for selecting the most informative visual tokens. It calculates both text-to-image similarity and visual saliency extracted from the vision encoder. By combining these two indicators, AdaFV identifies visual tokens that are not only semantically aligned with the text but also visually prominent or significant. TRIM (Song et al., 2025a) introduces a unique method that begins by identifying outlier tokens based on the similarity between text and visual tokens; these outliers are deemed important. Subsequently, a clustering algorithm is utilized to merge the remaining, less critical tokens. This approach prioritizes distinct, highly relevant tokens before consolidating the rest.
3.4.3 Analysis of Similarity Methods
While query-based methods can precisely retain query-relevant tokens compared to the three prior approaches, they are unsuitable for multi-turn QA scenarios. This is because the initial query’s token retention is based on its specific question. A subsequent query may target different tokens, necessitating a re-execution of the token compression process. This makes the approach highly inefficient for multi-turn conversations.
Table 3: Comparison of Training-Free Token Compression Methods for Video LLMs in Understanding Tasks